OpenStudy (anonymous):

State the domain of each function. g(x) = sqrt (x^2 - 9) I'm a little confused. Do I just need to plug in random numbers?

OpenStudy (anonymous):

no graph it first https://www.desmos.com/calculator

OpenStudy (jdoe0001):

the nick seems to say, " I like chicken fever" not sure if that's something to look forward to, but who knows

OpenStudy (anonymous):

domain are the x values

OpenStudy (anonymous):

x values that are defined :)

OpenStudy (anonymous):

domain is all real numbers except -3<x>3

OpenStudy (anonymous):

domain is : \[x^2-9 \ge0 \rightarrow |x|\ge3 \rightarrow x \in(-\infty,-3]\cup[3,\infty)\]

OpenStudy (anonymous):

or that

OpenStudy (anonymous):

the range is usually found by finding the domain of the inverse, however , in this question you can notice that the minimum occurs at x=0 which is -9 , and THE VALUE of function increases no matter where x goes , that is , the range is: \[[-9,\infty)\]

OpenStudy (anonymous):

Thank you.